OpenStudy (anonymous):

Two towns that are 31.5 km apart are 3 cm apart on a map. What is the scale of the map?

OpenStudy (studygurl14):

the team played how many games?

OpenStudy (anonymous):

the team played 42 games

OpenStudy (studygurl14):

use proportions: \[\frac{4}{12} = \frac{W}{42}\] Solve for W

OpenStudy (anonymous):

would W equal 3 then?

OpenStudy (anonymous):

3 is the proportional because there both dividable by 3?

OpenStudy (anonymous):

do u know how to do it still?

OpenStudy (studygurl14):

No, times 42 by 4, and 12 by W to get 42(4) = 12W

OpenStudy (anonymous):

no the answer is 14 u divid 42/12 and times 4 and u get 14

OpenStudy (studygurl14):

right!

OpenStudy (justnick09):

\[3/7=.42857, 1/6=.16666 \]

OpenStudy (justnick09):

now muiltiply both of those by 1092

OpenStudy (justnick09):

\[1092-182-468=442\]

OpenStudy (anonymous):

then the answer i get will be my answer

OpenStudy (justnick09):

yes ur answer is 442

OpenStudy (anonymous):

thnx!!!

OpenStudy (justnick09):

np medal and fan meh nd check out my youtube cover please

OpenStudy (anonymous):

i did fan u and wats ur youtube name

OpenStudy (justnick09):

^^ link in my question

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

i have another question?

OpenStudy (justnick09):

shoot

OpenStudy (anonymous):

lol its ok if u dont want to help?

OpenStudy (anonymous):

can u help me?????

OpenStudy (anonymous):

plezzzz